# Attainability of quantum state discrimination bounds with collective measurements on finite copies

### v1

@@ Abstract
One of the fundamental tenets of quantum mechanics is that non-orthogonal states cannot be distinguished perfectly. When distinguishing multiple copies of a mixed quantum state, a collective measurement, which generates entanglement between the different copies of the unknown state, can achieve a lower error probability than non-entangling measurements. The error probability that can be attained using a collective measurement on a finite number of copies of the unknown state is given by the Helstrom bound. In the limit where we can perform a collective measurement on asymptotically many copies of the quantum state, the quantum Chernoff bound gives the attainable error probability. It is natural to ask at what rate does the error tend to this asymptotic limit, and whether the asymptotic limit can be attained for any finite number of copies. In this paper we address these questions. We find analytic expressions for the Helstrom bound for arbitrarily many copies of the unknown state in several simple qubit examples. Using these analytic expressions, we investigate how the attainable error rate changes as we allow collective measurements on finite numbers of copies of the quantum state. We also investigate the necessary conditions to saturate the $M$-copy Helstrom bound. It is known that a collective measurement on all $M$-copies of the unknown state is always sufficient to saturate the $M$-copy Helstrom bound. However, general conditions for when such a measurement is necessary to saturate the Helstrom bound remain unknown. We investigate specific measurement strategies which involve entangling operations on fewer than all $M$-copies of the unknown state. For many regimes we find that a collective measurement on all $M$-copies of the unknown state is necessary to saturate the $M$-copy Helstrom bound.
@@ I Introduction
Quantum state discrimination can be described as the task of discriminating between $M$ copies of two unknown quantum states $\rho_{+}^{\otimes M}$ or $\rho_{-}^{\otimes M}$. Due to the no-cloning theorem, it is impossible to perform this task perfectly unless the states to be discriminated are orthogonal to one another wootters1982single . Therefore, it has become an important task to find measurements which minimise the average probability of making an error while remaining experimentally feasible. When there is only a single copy of the unknown quantum state, i.e. $M=1$, the optimal measurement is simple to implement, not requiring any entangling operations. Along this line, there have been several experimental studies focused on optimally distinguishing single copies of pure states cook2007optical ; wittmann2008demonstration ; bartuuvskova2008programmable ; waldherr2012distinguishing ; becerra2013implementation ; izumi2020experimental ; izumi2021adaptive ; sidhu2021quantum ; gomez2022experimental . For $M>1$, there are two main classes of measurements: (1) adaptive measurements which rely only on local operations and classical communication (LOCC) between the $M$ different modes of the unknown quantum states and (2) collective measurements which generate entanglement between all $M$ copies of the unknown quantum state. Clearly the latter is a more powerful class of measurement which includes the former as a special case. However, the cost of this is that collective measurements are more experimentally complex to implement. As a result of this, adaptive measurements, or LOCC measurements, have been well studied and shown to outperform non-adaptive measurements walgate2000local ; becerra2013experimental ; higgins2009mixed . In contrast, the advantage of collective measurements over LOCC measurements has only been demonstrated recently conlon2023discriminating ; zhou2023experimental . Note that the advantage of collective measurements has been demonstrated in several other areas of quantum information roccia2017entangling ; hou2018deterministic ; conlon2023approaching ; parniak2018beating ; wu2019experimentally ; yuan2020direct ; wu2020minimizing . Throughout this work we shall refer to collective measurements as those acting on multiple copies of the same state [^1] However, it is important to note that many other tasks in quantum information also benefit from measurements in an entangled basis, which do not necessarily act on multiple copies of the same quantum state, such as quantum metrology marciniak2022optimal , communications delaney2022demonstration ; crossman2023quantum and orienteering gisin1999spin ; jeffrey2006optical ; tang2020experimental ..
@@
Theoretically, the role of collective measurements in state discrimination is reasonably well understood. Helstrom found a simple expression for the minimal error probability that can be attained when performing an entangling measurement on all $M$-copies of the unknown state. This has become known as the Helstrom bound helstrom1969quantum . Helstrom’s work also provides the entangling measurement which can saturate his bound. In certain situations however, collective measurements are not needed to saturate the Helstrom bound. Trivially, when $M=1$, there is only one copy of the unknown state available and so collective measurements are not possible. When discriminating two pure states it is known that collective measurements offer no advantage over separable measurements brody1996minimum ; ban1997accessible ; acin2005multiple . (Note however, that in slightly different settings collective measurements may outperform separable measurements for discriminating certain families of pure states peres1991optimal ; massar1995optimal ; chitambar2013revisiting .) Additionally, when the two states to be discriminated are simultaneously diagonalisable, the Helstrom bound can be saturated through the maximum likelihood measurement which requires only local operations, see Ref. audenaert2014upper and appendix A. However, in general for $M>1$ when discriminating between two mixed states $\rho_{+}^{\otimes M}$ or $\rho_{-}^{\otimes M}$, collective measurements may be required to saturate the Helstrom bound. In spite of the importance of this question, concrete results regarding when collective measurements are required to saturate the Helstrom bound remain largely unknown (Necessary and sufficient conditions that the optimal measurement must satisfy are known, see e.g. Refs. helstrom1969quantum ; ban1997optimum and appendix B). Much of our knowledge in this area comes from numerical results, where it has been shown that for discriminating qubit states collective measurements outperform separable measurements calsamiglia2010local ; higgins2011multiple ; conlon2023discriminating .
@@
When we take the limit of performing collective measurements on a large number of copies of the unknown state, the error probability is given by the quantum Chernoff bound audenaert2007discriminating ; Nussbaum2009chernoff . This is the quantum analogue of the Chernoff bound in classical statistics chernoff1952measure . Interestingly, the conditions for what measurement strategy is optimal changes in the asymptotic limit. For example, it is known that a fixed single-copy measurement is capable of asymptotically approaching this minimum error probability when discriminating two states, one of which is pure acin2005multiple ; calsamiglia2008quantum . Additionally, we note that Hayashi has shown that LOCC measurements do not outperform fixed local measurements in terms of the asymptotic error rate hayashi2009discrimination . Tables 1 and 2 summarise these known results regarding when different measurement strategies and bounds are equal. We compare both the asymptotic (Table 1) and finite-copy (Table 2) regime. These tables also indicate where we see our work as fitting into the existing literature.
@@
In this paper we investigate the attainability of both the quantum Chernoff bound and the Helstrom bound given finite resources. For the attainability of the quantum Chernoff bound we investigate the rate at which the error probability specified by the quantum Chernoff bound is approached with collective measurements on a finite number of copies of the unknown state. We do this by comparing the error probability provided by the quantum Chernoff bound to that of the Helstrom bound. We also investigate whether or not the quantum Chernoff bound is attainable given any finite number of copies of the unknown state. Essentially this amounts to asking whether the error probability provided by the Helstrom bound (when appropriately normalised for the number of copies of the unknown state being considered) ever reaches the error probability specified by the quantum Chernoff bound.
@@
For investigating the attainability of the Helstrom bound we ask whether collective measurements on all $M$ copies of a quantum state are necessary to saturate the Helstrom bound. We address this question using parameterised quantum circuits optimised to find measurements which minimise the error probablity. Although there exists much numerical evidence that there is a gap between the Helstrom bound and the LOCC bound Ref. calsamiglia2010local , to the best of our knowledge whether the Helstrom bound can be saturated by intermediate measurements remains unknown and relatively unexplored. By intermediate measurements we mean measurements capable of generating entanglement between more than 1, but not all $M$ copies of the unknown quantum state. We investigate how close these intermediate measurements can be to the Helstrom bound. Our investigations also reveal the optimal strategies given limited entangling resources. For example, given $M$ copies of the unknown state, and a device capable of faithfully performing an entangling measurement on $90\%$ of these copies, is it better to perform this entangling operation on the first $90\%$ or the last $90\%$? Our results indicate that the later option is superior.
@@
We note that there have been several previous studies which have examined state discrimination in the finite copy setting. Refs. pereira2023analytical ; audenaert2012quantum ; rouze2017finite ; li2014second provides bounds on the error probability attainable in the $M$-copy asymmetric state discrimination setting. As we provide exact expressions for the $M$-copy Helstrom bound, our work can be thought of as building on these previous more general works, by providing exact solutions for several simple examples. We also note that there are several works bounding other closely related quantities, including the channel capacity and quantum channel discrimination, in the finite-copy regime tomamichel2013hierarchy ; wang2012one ; tomamichel2016quantum ; bergh2024parallelization .
@@
The rest of the paper is set out as follows. In section II, we introduce the preliminary material: the Helstrom bound, the quantum Chernoff bound and the questions we wish to address. In section III, we present our main results on the attainability of the quantum Chernoff bound. We present analytic expressions for the Helstrom bound when discriminating multiple copies of qubit states. We use this to investigate the rate at which the Chernoff bound is approached for large numbers of copies of the unknown state. In section IV we investigate the requirements to saturate the $M$-copy Helstrom bound, building on the results in the previous section. Finally, in section V, we present our final remarks and conclude.
@@
$\kappa$ Equal by definition audenaert2007discriminating ; Nussbaum2009chernoff . Equal for pure states brody1996minimum ; ban1997accessible ; acin2005multiple . Equal up to a constant when one of the states is pure acin2005multiple ; calsamiglia2008quantum .
$P^{\text{H}}_{\text{e}}(M)$ Equal for pure states brody1996minimum ; ban1997accessible ; acin2005multiple .
Gap between the two conjectured to persist in the asymptotic limit for mixed qubit states in Ref. calsamiglia2010local .
By virtue of the previous column, conjectured to be unequal in the asymptotic limit for mixed qubit states calsamiglia2010local .
By virtue of the above row, equal up to a constant when one state is pure acin2005multiple ; calsamiglia2008quantum .
$P_{\text{e,LOCC}}$ By virtue of the above row, equal up to a constant when one state is pure acin2005multiple ; calsamiglia2008quantum .
Not equal for mixed states higgins2011multiple .
Fixed measurements can achieve the same scaling as adaptive hayashi2009discrimination .
Table 2: Conditions for equality between different measurement strategies in quantum state discrimination in the finite copy limit. Each box gives the known results about the conditions for equality between the corresponding measurement strategies or bounds in that row and column. $\kappa$, $P^{\text{H}}_{\text{e}}(M)$, $P_{\text{e,LOCC}}$ and $P_{\text{e,LO}}$ denote the $M$-copy quantum Chernoff bound, the $M$-copy Helstrom bound, the LOCC error probability and the error probability when performing the same local operation respectively. In this table we consider any finite positive integer value of $M$
@@
$P^{\text{H}}_{\text{e}}(M)$ For qubit states can be equal or unequal depending on the number of copies and the prior probability conlon2023discriminating ; calsamiglia2010local .
Equal for pure states brody1996minimum ; ban1997accessible ; acin2005multiple .
@@
In general not equal higgins2011multiple ; calsamiglia2010local .
Equal for any two simultaneously diagonalisable states through the maximum likelihood measurement, see Ref. audenaert2014upper and appendix A.
$P_{\text{e,LOCC}}$ In general not equal, see for example Ref higgins2011multiple ; calsamiglia2010local
@@ II.1 Helstrom bound
In general, the optimal measurement to distinguish $\rho_{\pm}^{\otimes M}$, will be a collective measurement which involves entangling operations between all $M$ modes. The minimum error probability, in this case, is given by the Helstrom bound helstrom1969quantum
@@
The state $\rho_{\pm}$ is guessed according to the sign of the measurement output. Note that, in some scenarios, $\Gamma$ will be either positive or negative meaning that the best strategy is always simply to guess the corresponding state. In practice, we implement the operator $\Gamma$, through a positive operator valued measure (POVM). A POVM is a set of positive operators $\{\Pi_{k}\geq 0\}$, which sum to the identity $\sum_{k}\Pi_{k}=\mathbb{I}$. Using the optimal measurement, given in Eq. (3), and the definition of error probability in Eq. (1), we can verify that the Helstrom bound becomes $(1-\lVert\Gamma\rVert)/2$, as expected from Eq. (2bergou2010discrimination . In general the POVM required to implement $\Gamma$ will correspond to a collective measurement.
As we wish to compare errors when distinguishing different numbers of quantum states, it is important to normalise for the number of states being considered appropriately. Otherwise, given arbitrarily many copies of the state one can still get an arbitrarily small error probability by measuring the unknown states individually. To determine the appropriate method for normalisation we shall consider the quantum Chernoff bound
@@ II.2 Quantum Chernoff bound
The quantum Chernoff bound (hereafter abbreviated to Chernoff bound) gives the error probability when we allow for collective measurements on asymptotically many copies of the unknown state. Specifically audenaert2007discriminating ; Nussbaum2009chernoff
@@
and $P_{\text{e,min},M}$ is the minimum error probability given $M$ copies of the probe state, i.e. $P_{\text{e,min},M}=P^{\text{H}}_{\text{e}}(M)$.
Noting that
$\lim_{M\to\infty}P_{\text{e,min},M}^{1/M}=\kappa\;.$ (7)
Note that for distinguishing more than two states, the Chernoff bound can be off by a constant factor audenaert2014upper . However, for distinguishing two states this effect is not present. By comparing $\epsilon_{M}$ and $\epsilon_{\infty}$ we can examine how rapidly the ultimate limits on the error exponent can be approached.
When distinguishing quantum states we find situations where $\epsilon_{1}\geq\epsilon_{2}\geq\epsilon_{3}\ldots\geq\epsilon_{\infty}$, or equivalently $P_{\text{e}}(1)\leq P_{\text{e}}(2)^{1/2}\leq P_{\text{e}}(3)^{1/3}\ldots\leq\kappa$. At first glance this may appear surprising, however this is entirely in keeping with the results of Refs. audenaert2007discriminating ; Nussbaum2009chernoff . Specifically Ref. Nussbaum2009chernoff proved
and Ref. audenaert2007discriminating derived
which implies
$\frac{\text{log}(P_{\text{e,min},M})}{M}\leq\text{log}(\kappa)\;.$ (11)
II.3 Quantities to investigate the error scaling
We are now well positioned to investigate various questions related to the attainability of the Chernoff bound. First we wish to know whether the error probability given by the Chernoff bound is attainable with any finite number of copies of the unknown state. To clarify what we mean by this, we have plotted two functions in Fig. 1. For one of these functions, the asymptotic limit of the function is attainable for finite arguments, for the other it is not. As the Chernoff bound specifically applies only in the asymptotic limit, it may be expected that the asymptotic limit cannot be reached for any finite argument, as is the case for many mathematical functions. Nevertheless, it is still worthwhile to investigate this, and related, questions. For instance, when investigating this question, we found interesting results about the rate at which the Chernoff bound is approached.
Let us define
$\delta(M,N)=R(M)-R(N)\;,$ (13)
Let us also define
$R_{\epsilon}(M)=\frac{\epsilon’_{M}}{\epsilon_{\infty}}\;,$ (14)
and
$\delta_{\epsilon}(M,N)=\frac{1}{\epsilon_{\infty}}(\epsilon’_{M}-\epsilon’_{N})\;,$ (15)
where $N>M$ and
$\epsilon’_{M}=-\frac{\text{log}(P_{\text{e,min},M})-\text{log}(q^{s}(1-q)^{1-s})}{M}.$
These quantities provide a similar comparison in terms of the error exponent. Although, this provides essentially the same information as the above quantities, it is not as physically well motivated. This is because the Helstrom bound gives the error probability when a finite number of copies are considered, which cannot be used to directly obtain the error exponent when the number of copies is small. To get around this, later in the manuscript, we consider the error exponent attained when repeatedly performing the Helstrom measurement.

### v2

@@ Abstract
One of the fundamental tenets of quantum mechanics is that non-orthogonal states cannot be distinguished perfectly. When distinguishing multiple copies of a mixed quantum state, a collective measurement, which generates entanglement between the different copies of the unknown state, can achieve a lower error probability than non-entangling measurements. The error probability that can be attained using a collective measurement on a finite number of copies of the unknown state is given by the Helstrom bound. In the limit where we can perform a collective measurement on asymptotically many copies of the quantum state, the quantum Chernoff bound gives the attainable error probability. It is natural to ask then what strategies can be employed to reach these two bounds – is entanglement across all available modes always necessary or can other, experimentally more simple, measurements saturate these bounds. In this work we address this question. We find analytic expressions for the Helstrom bound for arbitrarily many copies of the unknown state in simple qubit examples. Using these analytic expressions, we investigate whether the quantum Chernoff bound can be saturated by repeatedly implementing the $M$-copy Helstrom measurement. We also investigate the necessary conditions to saturate the $M$-copy Helstrom bound. It is known that a collective measurement on all $M$-copies of the unknown state is always sufficient to saturate the $M$-copy Helstrom bound. However, general conditions for when such a measurement is necessary to saturate the Helstrom bound remain unknown. We investigate specific measurement strategies which involve entangling operations on fewer than all $M$-copies of the unknown state. For many regimes we find that a collective measurement on all $M$-copies of the unknown state is necessary to saturate the $M$-copy Helstrom bound.
@@ I Introduction
Quantum state discrimination can be described as the task of discriminating between $M$ copies of two unknown quantum states $\rho_{+}^{\otimes M}$ or $\rho_{-}^{\otimes M}$. Due to the no-cloning theorem, it is impossible to perform this task perfectly unless the states to be discriminated are orthogonal to one another [1, dieks1982communication]. Therefore, it has become an important task to find measurements which minimise the average probability of making an error while remaining experimentally feasible. When there is only a single copy of the unknown quantum state, i.e. $M=1$, the optimal measurement is simple to implement, not requiring any entangling operations. Along this line, there have been several experimental studies focused on optimally distinguishing single copies of pure states [2, 3, 4, 5, 6, 7, 8, 9, 10]. For $M>1$, there are two main classes of measurements: (1) adaptive measurements which rely only on local operations and classical communication (LOCC) between the $M$ different modes of the unknown quantum states and (2) collective measurements which generate entanglement between all $M$ copies of the unknown quantum state. Clearly the latter is a more powerful class of measurement which includes the former as a special case. However, the cost of this is that collective measurements are more experimentally complex to implement. As a result of this, adaptive measurements, or LOCC measurements, have been well studied and shown to outperform non-adaptive measurements [11, 12, 13]. In contrast, the advantage of collective measurements over LOCC measurements in this context has only been demonstrated experimentally recently [14, 15]. Note that the advantage of collective measurements has been experimentally demonstrated in several other areas of quantum information [16, 17, 18, 19, 20, 21, 22]. Throughout this work we shall refer to collective measurements as those acting on multiple copies of the same state [^1] However, it is important to note that many other tasks in quantum information also benefit from measurements in an entangled basis, which do not necessarily act on multiple copies of the same quantum state, such as quantum metrology [23], communications [24, 25] and orienteering [26, 27, 28]..
@@
Theoretically, the role of collective measurements in state discrimination is reasonably well understood. Helstrom found a simple expression for the minimal error probability that can be attained when performing an entangling measurement on all $M$-copies of the unknown state. This has become known as the Helstrom bound [29]. Helstrom’s work also provides the entangling measurement which can saturate his bound. In certain situations however, collective measurements are not needed to saturate the Helstrom bound. Trivially, when $M=1$, there is only one copy of the unknown state available and so collective measurements are not possible. When discriminating two pure states it is known that collective measurements offer no advantage over separable measurements [30, 31, 32]. (Note however, that in slightly different settings collective measurements may outperform separable measurements for discriminating certain families of pure states [33, 34, 35].) Additionally, when the two states to be discriminated are simultaneously diagonalisable, the Helstrom bound can be saturated through the maximum likelihood measurement which requires only local operations, see Ref. [36] and appendix A. However, in general for $M>1$ when discriminating between two mixed states $\rho_{+}^{\otimes M}$ or $\rho_{-}^{\otimes M}$, collective measurements may be required to saturate the Helstrom bound. In spite of the importance of this question, concrete results regarding when collective measurements are required to saturate the Helstrom bound remain largely unknown (Necessary and sufficient conditions that the optimal measurement must satisfy are known, see e.g. Refs. [29, 37] and appendix B). Much of our knowledge in this area comes from numerical results, where it has been shown that for discriminating qubit states collective measurements outperform separable measurements [38, 39, 14].
@@
When we take the limit of performing collective measurements on a large number of copies of the unknown state, the error probability is given by the quantum Chernoff bound [40, 41]. This is the quantum analogue of the Chernoff bound in classical statistics [42]. Interestingly, the conditions for what measurement strategy is optimal changes in the asymptotic limit. For example, it is known that a fixed single-copy measurement is capable of asymptotically approaching this minimum error probability when discriminating two states, one of which is pure [32, 43]. Additionally, we note that Hayashi has shown that LOCC measurements do not outperform fixed local measurements in terms of the asymptotic error rate [44]. Tables 1 and 2 summarise these known results regarding when different measurement strategies and bounds are equal. We compare both the asymptotic (Table 1) and finite-copy (Table 2) regime. These tables also indicate where we see our work as fitting into the existing literature.
@@
In this paper we investigate the attainability of both the quantum Chernoff bound and the Helstrom bound given finite resources. For the attainability of the quantum Chernoff bound we investigate the rate at which the error probability specified by the quantum Chernoff bound is approached with collective measurements on a finite number of copies of the unknown state. We do this by considering repeated implementations of the Helstrom measurement on groups of quantum states. This investigates the scenario where entangling measurements can be repeatedly implemented on small subsets of the available quantum states.
@@
For investigating the attainability of the Helstrom bound we ask whether collective measurements on all $M$ copies of a quantum state are necessary to saturate the Helstrom bound. We address this question using parameterised quantum circuits optimised to find measurements which minimise the error probablity. Although there exists much numerical evidence that there is a gap between the Helstrom bound and the LOCC bound Ref. [38], to the best of our knowledge whether the Helstrom bound can be saturated by intermediate measurements remains unknown and relatively unexplored. By intermediate measurements we mean measurements capable of generating entanglement between more than 1, but not all $M$ copies of the unknown quantum state. We investigate how close these intermediate measurements can be to the Helstrom bound. Our investigations also reveal the optimal strategies given limited entangling resources. For example, given $M$ copies of the unknown state, and a device capable of faithfully performing an entangling measurement on $90\%$ of these copies, is it better to perform this entangling operation on the first $90\%$ or the last $90\%$? Our results indicate that the later option is superior. We then extend these results from LOCC measurements to positive partial transpose (PPT) measurements which are a stronger class of measurements than LOCC, but still weaker than entangling measurements on all $M$ copies. Interestingly in this setting we find different results depending on whether we consider discriminating between two or more quantum states.
@@
We note that there have been several previous studies which have examined state discrimination in the finite copy setting. Refs. [45, 46, 47, 48] provides bounds on the error probability attainable in the $M$-copy asymmetric state discrimination setting. As we provide exact expressions for the $M$-copy Helstrom bound, our work can be thought of as building on these previous more general works, by providing exact solutions for several simple examples. We also note that there are several works bounding other closely related quantities, including the channel capacity and quantum channel discrimination, in the finite-copy regime [49, 50, 51, 52].
@@
The rest of the paper is set out as follows. In section II, we introduce the preliminary material: the Helstrom bound, the quantum Chernoff bound and the questions we wish to address. In section III, we present our main results on the attainability of the quantum Chernoff bound. We present analytic expressions for the Helstrom bound when discriminating multiple copies of qubit states. We use this to investigate the attainability of the Chernoff bound when repeatedly implementing the $M$-copy Helstrom measurement. In section IV we investigate the requirements to saturate the $M$-copy Helstrom bound, building on the results in the previous section. Finally, in section V, we present our final remarks and conclude.
@@
$\kappa$ Equal by definition [40, 41]. Equal for pure states [30, 31, 32]. Equal up to a constant when one of the states is pure [32, 43].
$P^{\text{H}}_{\text{e}}(M)$                Equal for pure states [30, 31, 32].
Gap between the two conjectured to persist in the asymptotic limit for mixed qubit states in Ref. [38].
By virtue of the previous column, conjectured to be unequal in the asymptotic limit for mixed qubit states [38].
By virtue of the above row, equal up to a constant when one state is pure [32, 43].
$P_{\text{e,LOCC}}$                               By virtue of the above row, equal up to a constant when one state is pure [32, 43].
Not equal for mixed states [39].
Fixed measurements can achieve the same scaling as adaptive [44].
Table 2: Conditions for equality between different measurement strategies in quantum state discrimination in the finite copy limit. Each box gives the known results about the conditions for equality between the corresponding measurement strategies or bounds in that row and column. $\kappa$, $P^{\text{H}}_{\text{e}}(M)$, $P_{\text{e,LOCC}}$ and $P_{\text{e,LO}}$ denote the $M$-copy quantum Chernoff bound, the $M$-copy Helstrom bound, the LOCC error probability and the error probability when performing the same local operation respectively. In this table we consider any finite positive integer value of $M$. Note that we include the quantum Chernoff bound although it is, by definition, an asymptotic limit. For this row, we refer to repeated implementations of the corresponding optimal $M$-copy measurement, and whether that can asymptotically saturate the quantum Chernoff bound.
@@
$P^{\text{H}}_{\text{e}}(M)$                For qubit states can be equal or unequal depending on the number of copies and the prior probability [14, 38].
Equal for pure states [30, 31, 32].
@@
In general not equal [39, 38].
Equal for any two simultaneously diagonalisable states through the maximum likelihood measurement, see Ref. [36] and appendix A.
$P_{\text{e,LOCC}}$                                   In general not equal, see for example Refs [39, 38]
@@ II.1 Helstrom bound
In general, the optimal measurement to distinguish $\rho_{\pm}^{\otimes M}$, will be a collective measurement which involves entangling operations between all $M$ modes. The minimum error probability, in this case, is given by the Helstrom bound [29]
@@
The state $\rho_{\pm}$ is guessed according to the sign of the measurement output. Note that, in some scenarios, $\Gamma$ will be either positive or negative meaning that the best strategy is always simply to guess the corresponding state. In practice, we implement the operator $\Gamma$, through a positive operator valued measure (POVM). A POVM is a set of positive operators $\{\Pi_{k}\geq 0\}$, which sum to the identity $\sum_{k}\Pi_{k}=\mathbb{I}$. Using the optimal measurement, given in Eq. (3), and the definition of error probability in Eq. (1), we can verify that the Helstrom bound becomes $(1-\lVert\Gamma\rVert)/2$, as expected from Eq. (2[53]. In general the POVM required to implement $\Gamma$ will correspond to a collective measurement.
@@ II.2 Quantum Chernoff bound
The quantum Chernoff bound (hereafter abbreviated to Chernoff bound) gives the error probability when we allow for collective measurements on asymptotically many copies of the unknown state. Specifically [40, 41]
@@
and $P_{\text{e,min},M}$ is the minimum error probability given $M$ copies of the probe state, i.e. $P_{\text{e,min},M}=P^{\text{H}}_{\text{e}}(M)$. Given $M\gg 1$ copies of a quantum state, we can expect the error probability to scale as $P_{\text{e,min},M}\propto\text{e}^{-\epsilon_{M}M}$. We call this $\epsilon_{M}$ term the error exponent.